OpenStudy (anonymous):

4. If a = b then a - c = b - c. What property is this? A. addition property of equality B. subtraction property of equality C. transitive property D. commutative property

OpenStudy (anonymous):

Transitive property. Because if a=b, then a-c = b-c.

OpenStudy (anonymous):

thank you i was having a hard time understanding

OpenStudy (anonymous):

5. Which statement below represents the transitive property? A. a = a B. If a = b then a + b = a + c C. If a = b then b = a. D. If a = b and b = c then a = c.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thank you thats what i thought

OpenStudy (anonymous):

Good thinking. Always remember that if two are equal and the third is equal to one of the two, it must be equal to the other one as well. Transitivity.

OpenStudy (anonymous):

ok thank you i will remember that